# On estimation and order selection for multivariate extremes via clustering

### v3

@@ 3.2.1 Some deterministic estimates
Fix throughout this section distinct points $\left\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\right\}=:A$ on $\mathbb{S}_{+}^{d-1}$, $k\in\mathbb{Z}_{+}$, $p_{i}>0$ with $p_{1}+\ldots+p_{k}=1$, and a dissimilarity measure $D$ defined in Definition 1. We say a multiset $W$ on $\mathbb{S}_{+}^{d-1}$ with $|W|<\infty$ satisfies the concentration condition $\mathcal{A}(\epsilon,\delta)$ with $\epsilon,\delta>0$, if $k\leq|W|\wedge\delta^{-1}$ such that $\epsilon\in(0,r_{A})$ with $r_{A}$ in (20), $\delta\in\left(0,p_{\min}\right)$ with $p_{\min}$ in (24), and
@@ 3.2.2 Proof of Theorem 1
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$ form an $m$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. If $m<k$, then almost surely,
@@
$E_{n}(\epsilon,\delta)=\{|W_{n}\cap B_{D}(\mathbf{a}_{i},\epsilon)|\geq|W_{n}|(p_{i}-\delta)\text{ for all }i\in\left\{1,\ldots,k\right\}\}.$ (34)
@@
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$ form an $m$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. Suppose $t>0$. If $m>k$, we have almost surely
@@
$\begin{cases}\limsup_{n}S_{t}(W_{n};A_{m,n},\mathfrak{C}_{m,n})\leq 1-r_{A}p_{\min},&\text{ if }m<k;\\ \liminf_{n}S_{t}(W_{n};A_{k,n},\mathfrak{C}_{k,n})\geq\left(r_{A}kp_{\min}\right)^{t};&\text{ if }m=k;\\ \limsup_{n}S_{t}(W_{n};A_{m,n},\mathfrak{C}_{m,n})\leq 0,&\text{ if }m>k.\end{cases}$
@@ 4 Large deviation analysis of clustering-based spectral estimation
We only prove the (35) and the proof of (36) is similar. It follows from a version of Hoeffding’s inequality for Binomial [18, Equation (2.1)] that for any $m\geq 0$,
@@
where in the last inequality we have used the inequality $x+1\leq\exp\left(x\right)$, $x\in\mathbb{R}$. To obtain the second inequality in (35), it suffices to note that in view of [18, Equation (2.3)] one has $\mathcal{D}(q_{1}-r|q_{1})\geq 2r^{2}$. ∎
@@
Suppose $\mathbf{X}$ satisfying (1) and (2) has a spectral measure of the following form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left({A}_{k,n}=\left({\mathbf{a}}_{1,n}^{k},\ldots,{\mathbf{a}}_{k,n}^{k}\right),\mathfrak{C}_{k,n}=\left\{{C}_{1,n}^{k},\ldots,{C}_{k,n}^{k}\right\}\right)$ form an $k$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. Let $E(x,y)$ be the event defined in 37. Then for any $x,y>0$,
@@
$\Delta(x,y)=\begin{cases}\max\{y/c_{k},p_{min}x/(k+x)\}&\text{when }x<\epsilon_{0}\text{, }y<c_{k}p_{\min}\epsilon_{0}/(k+\epsilon_{0}),\\ p_{\min}\epsilon_{0}/(k+\epsilon_{0})&otherwise,\end{cases}$
@@
$\Pr\left(\bigcup_{i=1}^{k}\{H_{n}\left(B_{D}(\mathbf{a}_{i},\epsilon)\right)<p_{i}-\delta\}\right)$ $\leq\sum_{i=1}^{k}\Pr\left(H_{n}(B_{D}(\mathbf{a}_{i},\epsilon))<p_{i}-\delta\right)$
$\leq k\exp\left(nq_{2}(n)\left[\exp\left\{-2\delta^{2}\right\}-1\right]\right).$ (40)
@@
$\begin{cases}\delta<p_{\min}\epsilon_{0}/(k+\epsilon_{0})&\text{if }x\geq\epsilon_{0},\\ \delta<p_{\min}\epsilon_{0}/(k+\epsilon_{0})&\text{if }x<\epsilon_{0},y\geq c_{k}p_{\min}\epsilon_{0}/(k+\epsilon_{0}),\\ \delta<\max\{y/c_{k},p_{min}x/(k+x)\}&\text{if }x<\epsilon_{0},y<c_{k}p_{\min}\epsilon_{0}/(k+\epsilon_{0}).\end{cases}$
@@
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, $k\in\mathbb{Z}_{+}$ where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$, $m\in\mathbb{Z}_{+}$, form an $m$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. Let $r_{A}$ be defined as in (20), $p_{\min}$ be defined as in (24) and $t_{0}$ be defined as in (25). Then fix $t\in(0,t_{0})$,
@@ 5.1 The models
As observed by Einmahl et al. [7] and Janßen and Wan [23], one may relate a $k$-clustering algorithm to the estimation of certain factor-like models that are often considered in the analysis of multivariate extremes. Suppose $B=\left(b_{ij}\right)_{i=1,\ldots,d,j=1,\ldots,k}=\left(\mathbf{b}_{1},\ldots,\mathbf{b}_{k}\right)$, where $\mathbf{b}_{j}=(b_{1j},\ldots,b_{dj})^{\top}$, $j\in\{1,\ldots,k\}$, are $k$ distinct $d$-dimensional vectors, $b_{ij}\geq 0$, and that each column and row vector of $B$ is nonzero (otherwise, the dimension $d$ or the factor order $k$ can be reduced). Assume that $\mathbf{Z}=(Z_{1},\ldots,Z_{k})^{\top}$ is a vector of i.i.d.  positive random variables satisfying $\Pr(Z_{1}>z)\sim z^{-\alpha}$ as $z\rightarrow\infty$, $\alpha\in(0,\infty)$. Then the sum-linear model is given as
@@
This can be derived based on the well-known “single large jump” heuristic: when $\|\mathbf{X}\|_{(r)}$ is large, it is only due to a single large $Z_{j}$ with overwhelming probability. See, e.g., [26] and [7]; we mention that these works usually assume the same norm $\|\cdot\|_{(r)}=\|\cdot\|_{(s)}$ and $\alpha=1$, although an extension is straightforward. In addition, the marginal standardization condition (1) or equivalently (7), imposes the following restriction on $B$:
@@ 6.1 Simulation studies
In this section, we present some simulation studies to illustrate the performance of the penalized ASW method introduced in Section 3. We follow the setup in [23, Section 4] to simulate the max-linear factor model (42) with randomly generated coefficient matrix $B$. In particular, we let the factors $Z_{j}$’s each follow a standard Fréchet ($\alpha=1$) distribution. We consider 4 different combinations of dimensionality $d$ and true order $k$. Under each $(d,k)$ combination, we describe in the list below the way the coefficient vector $\mathbf{b}_{j}$’s are generated. Note that due to the standardization (44), only $\mathbf{b}_{1},\ldots,\mathbf{b}_{k-1}$ need to be specified. Let $U_{i}$’s stand for i.i.d. uniform random variables on $[0,1]$.
@@
In Figures 2 $\sim$ 5, we demonstrate the simulation results through some graphical representations. Specifically, each colored matrix plot is associated with a $(d,k)$ setup as described above. In each plot, a column corresponds to a simulated dataset, and there are 100 columns. The upper half of the plot corresponds to spherical $k$-means and the lower half corresponds to $k$-pc. Within each of these halves, a row corresponds to a $t$ penalty parameter specification. The color of a cell in the matrix signifies the order $m$ chosen by maximizing the penalized ASW. We use a white color to indicate a coincidence of $m$ with the true order $k$, with a deeper shade of red indicating that the greater $m$ falls below the true $k$, and a deeper shade of blue indicating the greater it exceeds the true $k$. The bar graph to the right of the matrix indicates the success rate of order identification (that is, $m=k$) in all 100 instances.
@@
In the supplementary material [5], we include more simulation studies as above with different choices of extremal subsample sizes. The results are similar in terms of the bias correction effect of the penalty. We also observe that when the extremal subsample size is chosen relatively large (e.g., $1500$), the accuracy of order selection deteriorates, which is expected since the effect of the convergence of Corollary 1 is not yet strong with a low threshold.
Fig. 2: Simulation result visualization for the setup $d=4,k=2$ in Section 6.1.
Fig. 3: Simulation result visualization for the setup $d=4,k=6$ in Section 6.1.
Fig. 4: Simulation result visualization for the setup $d=6,k=6$ in Section 6.1.
Fig. 5: Simulation result visualization for the setup $d=10,k=6$ in Section 6.1.
@@ 6.2.1 Air Pollution Data
In Figures 7 and 9, following the same manner as in Figure 1, the penalized ASW is plotted against the number of clusters, where different curves correspond to different values of the tuning parameter $t$. With the visual method described in Section 3, we can identify orders as $5$ (although $4$ seems to be a reasonable choice as well) for the summer data and $3$ for the winter data respectively. These orders are similar to the choices $5$ for the summer data and $4$ for the winter data made in [23] under the guidance of certain elbow plots (see [23, Figure 1]). From the elbow plot in [23, Figure 1], it seems that $k=3$ for the winter data is also plausible. Recall also that here we use the spherical $k$-pc algorithm of [10] while [23] used the spherical $k$-means.
Following the method introduced in Section 5.2 with $\|\cdot\|_{(s)}=\|\cdot\|_{(r)}=\|\cdot\|_{2}$ and $\alpha=2$, we compute the factor coefficient matrix $B$ for the two datasets; see Figures 7 and 9. Furthermore, Figures 7 and 9 include visualizations of cluster centers computed based on the $k$-pc algorithm of [10] for the two datasets when we choose the numbers of clusters as above, respectively. Each row in either of the plots corresponds to the coordinate vector of a cluster center: a deeper shade of color indicates a higher value of the coordinate. For the summer data in Figure 7, whose order has been chosen as $5$, the cluster centers concentrate sharply near coordinate directions, which to an extent indicates an asymptotic (or say extremal) independence (see, e.g., [1, Chapter 8]) of the pollutants. In contrast, for the winter data in Figure 9, whose order has been chosen as $3$, a cluster center indicates a group of concomitant extremes consisting of NO, NO2 and PM10. The asymptotic dependence between these 3 variables has been observed in [17]. This serves as a support for our order choice which has placed these 3 variables in the same concomitant group.

### v4

@@ 3.2.1 Some deterministic estimates
Fix throughout this section distinct points $\left\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\right\}=:A$ on $\mathbb{S}_{+}^{d-1}$, $k\in\mathbb{Z}_{+}$, $p_{i}>0$ with $p_{1}+\cdots+p_{k}=1$, and a dissimilarity measure $D$ defined in Definition 1. We say a multiset $W$ on $\mathbb{S}_{+}^{d-1}$ with $|W|<\infty$ satisfies the concentration condition $\mathcal{A}(\epsilon,\delta)$ with $\epsilon,\delta>0$, if $k\leq|W|\wedge\delta^{-1}$ such that $\epsilon\in(0,r_{A})$ with $r_{A}$ in (20), $\delta\in\left(0,p_{\min}\right)$ with $p_{\min}$ in (24), and
@@ 3.2.2 Proof of Theorem 1
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\cdots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$ form an $m$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. If $m<k$, then almost surely,
@@
$E_{n}(\epsilon,\delta)=\left\{|W_{n}\cap B_{D}(\mathbf{a}_{i},\epsilon)|\geq|W_{n}|(p_{i}-\delta),\ i\in\left\{1,\ldots,k\right\}\right\}.$ (34)
@@
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\cdots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$ form an $m$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. Suppose $t>0$. If $m>k$, we have almost surely
@@
$\begin{cases}\limsup_{n}S_{t}(W_{n};A_{m,n},\mathfrak{C}_{m,n})\leq 1-r_{A}p_{\min},&\text{ if }m<k;\\ \liminf_{n}S_{t}(W_{n};A_{k,n},\mathfrak{C}_{k,n})\geq\left(r_{A}kp_{\min}\right)^{t},&\text{ if }m=k;\\ \limsup_{n}S_{t}(W_{n};A_{m,n},\mathfrak{C}_{m,n})\leq 0,&\text{ if }m>k.\end{cases}$
@@ 4 Large deviation analysis of clustering-based spectral estimation
We only prove the (35) and the proof of (36) is similar. It follows from a version of Hoeffding’s inequality for Binomial [17, Equation (2.1)] that for any $m\geq 0$,
@@
where in the last inequality we have used the inequality $x+1\leq\exp\left(x\right)$, $x\in\mathbb{R}$. To obtain the second inequality in (35), it suffices to note that in view of [17, Equation (2.3)] one has $\mathcal{D}(q_{1}-r|q_{1})\geq 2r^{2}$. ∎
@@
Suppose $\mathbf{X}$ satisfying (1) and (2) has a spectral measure of the following form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\cdots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left({A}_{k,n}=\left({\mathbf{a}}_{1,n}^{k},\ldots,{\mathbf{a}}_{k,n}^{k}\right),\mathfrak{C}_{k,n}=\left\{{C}_{1,n}^{k},\ldots,{C}_{k,n}^{k}\right\}\right)$ form an $k$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. Let $E(x,y)$ be the event defined in 37. Then for any $x,y>0$,
@@
$\Delta(x,y)=\begin{cases}\max\{y/c_{k},p_{min}x/(k+x)\},&x<\epsilon_{0}\text{,}y<c_{k}p_{\min}\epsilon_{0}/(k+\epsilon_{0}),\\ p_{\min}\epsilon_{0}/(k+\epsilon_{0}),&otherwise,\end{cases}$
@@
$\Pr\left(\bigcup_{i=1}^{k}\{H_{n}\left(B_{D}(\mathbf{a}_{i},\epsilon)\right)<p_{i}-\delta\}\right)$ $\leq\sum_{i=1}^{k}\Pr\left(H_{n}(B_{D}(\mathbf{a}_{i},\epsilon))<p_{i}-\delta\right)\leq k\exp\left(nq_{2}(n)\left[\exp\left\{-2\delta^{2}\right\}-1\right]\right).$ (40)
@@
$\begin{cases}\delta<p_{\min}\epsilon_{0}/(k+\epsilon_{0}),&\text{if }x\geq\epsilon_{0},\\ \delta<p_{\min}\epsilon_{0}/(k+\epsilon_{0}),&\text{if }x<\epsilon_{0},y\geq c_{k}p_{\min}\epsilon_{0}/(k+\epsilon_{0}),\\ \delta<\max\{y/c_{k},p_{min}x/(k+x)\},&\text{if }x<\epsilon_{0},y<c_{k}p_{\min}\epsilon_{0}/(k+\epsilon_{0}).\end{cases}$
@@
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, $k\in\mathbb{Z}_{+}$ where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\cdots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$, $m\in\mathbb{Z}_{+}$, form an $m$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. Let $r_{A}$ be defined as in (20), $p_{\min}$ be defined as in (24) and $t_{0}$ be defined as in (25). Then fix $t\in(0,t_{0})$,
@@ 5.1 The models
As observed by Einmahl et al. [6] and Janßen and Wan [22], one may relate a $k$-clustering algorithm to the estimation of certain factor-like models that are often considered in the analysis of multivariate extremes. Suppose $B=\left(b_{ij}\right)_{i\in\{1,\ldots,d\},j\in\{1,\ldots,k\}}=\left(\mathbf{b}_{1},\ldots,\mathbf{b}_{k}\right)$, where $\mathbf{b}_{j}=(b_{1j},\ldots,b_{dj})^{\top}$, $j\in\{1,\ldots,k\}$, are $k$ distinct $d$-dimensional vectors, $b_{ij}\geq 0$, and that each column and row vector of $B$ is nonzero (otherwise, the dimension $d$ or the factor order $k$ can be reduced). Assume that $\mathbf{Z}=(Z_{1},\ldots,Z_{k})^{\top}$ is a vector of i.i.d.  positive random variables satisfying $\Pr(Z_{1}>z)\sim z^{-\alpha}$ as $z\rightarrow\infty$, $\alpha\in(0,\infty)$. Then the sum-linear model is given as
@@
This can be derived based on the well-known “single large jump” heuristic: when $\|\mathbf{X}\|_{(r)}$ is large, it is only due to a single large $Z_{j}$ with overwhelming probability. See, e.g., [medina2024spectral, 6]; we mention that these works usually assume the same norm $\|\cdot\|_{(r)}=\|\cdot\|_{(s)}$ and $\alpha=1$, although an extension is straightforward. In addition, the marginal standardization condition (1) or equivalently (7), imposes the following restriction on $B$:
@@ 6.1 Simulation studies
In this section, we present some simulation studies to illustrate the performance of the penalized ASW method introduced in Section 3. We follow the setup in [22, Section 4] to simulate the max-linear factor model (42) with randomly generated coefficient matrix $B$. In particular, we let the factors $Z_{j}$’s each follow a standard Fréchet ($\alpha=1$) distribution. We consider 4 different combinations of dimensionality $d$ and true order $k$. Under each $(d,k)$ combination, we describe in the list below the way the coefficient vector $\mathbf{b}_{j}$’s are generated. Note that due to the standardization (44), only $\mathbf{b}_{1},\ldots,\mathbf{b}_{k-1}$ need to be specified. Let $U_{i}$’s stand for i.i.d. uniform random variables on $[0,1]$;
@@
In Fig. 2 $\sim$ 5, we demonstrate the simulation results through some graphical representations. Specifically, each colored matrix plot is associated with a $(d,k)$ setup as described above. In each plot, a column corresponds to a simulated dataset, and there are 100 columns. The upper half of the plot corresponds to spherical $k$-means and the lower half corresponds to $k$-pc. Within each of these halves, a row corresponds to a $t$ penalty parameter specification. The color of a cell in the matrix signifies the order $m$ chosen by maximizing the penalized ASW. We use a white color to indicate a coincidence of $m$ with the true order $k$, with a deeper shade of red indicating that the greater $m$ falls below the true $k$, and a deeper shade of blue indicating the greater it exceeds the true $k$. The bar graph to the right of the matrix indicates the success rate of order identification (that is, $m=k$) in all 100 instances.
@@
In the supplementary material, we include more simulation studies as above with different choices of extremal subsample sizes. The results are similar in terms of the bias correction effect of the penalty. We also observe that when the extremal subsample size is chosen relatively large (e.g., $1500$), the accuracy of order selection deteriorates, which is expected since the effect of the convergence of Corollary 1 is not yet strong with a low threshold.
Fig. 2: Simulation result visualization for the setup $d=4,k=2$ in Section 6.1. A column corresponds to a simulated dataset, and a row corresponds to a $t$ penalty parameter specification. See Section 6.1 for more details.
Fig. 3: Simulation result visualization for the setup $d=4,k=6$ in Section 6.1. A column corresponds to a simulated dataset, and a row corresponds to a $t$ penalty parameter specification. See Section 6.1 for more details.
Fig. 4: Simulation result visualization for the setup $d=6,k=6$ in Section 6.1. A column corresponds to a simulated dataset, and a row corresponds to a $t$ penalty parameter specification. See Section 6.1 for more details.
Fig. 5: Simulation result visualization for the setup $d=10,k=6$ in Section 6.1. A column corresponds to a simulated dataset, and a row corresponds to a $t$ penalty parameter specification. See Section 6.1 for more details.
@@ 6.2.1 Air Pollution Data
In Fig. 7 and 9, following the same manner as in Fig. 1, the penalized ASW is plotted against the number of clusters, where different curves correspond to different values of the tuning parameter $t$. With the visual method described in Section 3, we can identify orders as $5$ (although $4$ seems to be a reasonable choice as well) for the summer data and $3$ for the winter data respectively. These orders are similar to the choices $5$ for the summer data and $4$ for the winter data made in [22] under the guidance of certain elbow plots (see [22, Fig. 1]). From the elbow plot in [22, Fig. 1], it seems that $k=3$ for the winter data is also plausible. Recall also that here we use the spherical $k$-pc algorithm of [9] while [22] used the spherical $k$-means.
Following the method introduced in Section 5.2 with $\|\cdot\|_{(s)}=\|\cdot\|_{(r)}=\|\cdot\|_{2}$ and $\alpha=2$, we compute the factor coefficient matrix $B$ for the two datasets; see Fig. 7 and 9. For the summer data in Fig. 7, whose order has been chosen as $5$, the factor coordinates concentrate sharply near coordinate directions, which to an extent indicates an asymptotic (or say extremal) independence (see, e.g., [1, Chapter 8]) of the pollutants. In contrast, for the winter data in Fig. 9, whose order has been chosen as $3$, a factor indicates a group of concomitant extremes consisting of NO, NO2 and PM10. The asymptotic dependence between these 3 variables has been observed in [16]. This serves as a support for our order choice which has placed these 3 variables in the same concomitant group.